OpenStudy (tatumlee):

√x-3=8 can someone show me how to do this problem? thank you (:

OpenStudy (anonymous):

u want to find the value of x

OpenStudy (anonymous):

This is what I got Let's solve your equation step-by-step. √x−3=8 Step 1: Add 3 to both sides. √x−3+3=8+3 √x=11 Step 2: Solve Square Root. √x=11 x=112(Square both sides) x=121 Check answers. (Plug them in to make sure they work.) x=121(Works in original equation) Answer: x=121

OpenStudy (anonymous):

There you go

OpenStudy (anonymous):

isn sqrt on all x-3

OpenStudy (tatumlee):

Thank you so much that's really helpful, its actually simpler than I thought.

OpenStudy (anonymous):

I knew you had it in you